OpenStudy (gabylovesyou):

@misty1212 What are the zero(s) of the function f(x) = the quantity of 4 x squared minus 36 x, all over x minus 9? x = −9 x = 0 x = 9 x = 0 and x = 9

OpenStudy (gabylovesyou):

\[f(x) = \frac{ 4x^2-36x }{ x - 9 }\]

OpenStudy (gabylovesyou):

@satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (misty1212):

HI let me get rid of this bozo hold on

OpenStudy (anonymous):

as flapperreaper

OpenStudy (gabylovesyou):

@misty1212 lol okay.. and my answer to this is x = 0 ..

OpenStudy (misty1212):

\[f(x) = \frac{ 4x^2-36x }{ x - 9 }\] \[f(x) = \frac{ 4x(x-9) }{ x - 9 }=4x\]

OpenStudy (misty1212):

you are right, zero is at \(0\)

OpenStudy (gabylovesyou):

ok yay thanks :) ..

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]